ganeshie8 (ganeshie8):

Find the number of possible 4x4 matrices such that : 1) each row has two 0's and two 1's 2) each column has two 0's and two 1's 3) sum of elements in each row = sum of elements in each column = 2

ganeshie8 (ganeshie8):

example : \[\large \begin{pmatrix} 1&1&0&0\\1&1&0&0\\0&0&1&1\\0&0&1&1\end{pmatrix}\]

ganeshie8 (ganeshie8):

i think the 3rd constraint is redundant... but you get the idea ? ikram thinking of trying it in matlab ?

OpenStudy (dan815):

the problem seemed more cool when its called the zero sum game =[

OpenStudy (ikram002p):

hehe sort of , im trying with my friend :P

OpenStudy (dan815):

IRL friend?

OpenStudy (ikram002p):

IRL ?

OpenStudy (ikram002p):

ok my friend gave me this idea , \[\left(\begin{matrix}4 \\ 2\end{matrix}\right)^2 \]

OpenStudy (dan815):

tell her i hate her

ganeshie8 (ganeshie8):

okay have fun then :) it wont be easy to code in matlab

OpenStudy (ikram002p):

she is smarter than u :P

ganeshie8 (ganeshie8):

your friend seem to be having a good start... but thats not the answer

OpenStudy (dan815):

hey i found an interesting thing

OpenStudy (dan815):